# Cosmological Inflation and Dark Sector from 11D Supergravity

### v2

@@ Abstract
We explore compactifications of the form $\mathcal{M}^{3,1}\times T_{q_{1}}^{2}\times T_{q_{2}}^{2}\times T_{q_{3}}^{2}\times S^{1}$ in the framework of 11D supergravity. By imposing suitable gauge conditions and boundary conditions, we find that the four-dimensional FRW universe emerges as a solution representing cosmological D3-branes in the eleven-dimensional bulk. These specific compactification methods can produce cosmological inflation that aligns with the observational constraints set by the 2021 BICEP/Keck BICEP:2021xfz and Planck 2018 results Planck:2018jri . In the cosmological inflation models we construct, the inflaton can be interpreted as the conformal vibrations of extra dimensions with a size around $10^{5}$ times the reduced Planck length. Additionally, we offer the theoretical predictions for the mass of the inflaton, and the tree-level Newton’s gravity law between two massive point particles surrounded by a spherically symmetric distribution of the inflaton, which can reproduce the Tully-Fisher relation and explain the flat rotation curves of galaxies.
1 Introduction
@@
The eleven-dimensional supergravity is widely believed to be the tree-level low-energy effective theory of M-theory. As an ultraviolet complete theory, M-theory possesses inherent advantages in addressing extremely high-energy scenarios, such as the inflationary period of the early universe. However, our understanding of M-theory remains limited, see Banks:1996vh ; Polchinski:1999br ; Taylor:2001vb for remarkable progress in M-theory. A comprehensive theoretical description is still necessary for understanding physics in the extremely early universe by M-theory. From the perspective of string theory, it is conjectured that by compactifying M-theory on a special manifold of $G_{2}$ holonomy, it could yield the supersymmetric version of Standard Model of Particle Physics and General Relativity Acharya:2007rc ; Kane:2011kj . This conjecture has generated some discussions on M-theory cosmology in recent years, particularly regarding $G_{2}$ compactification, see Kane:2019nod ; Brustein:2002xf ; Coley:2002xb for examples. While this paper does not aim to fully realize the ambitious goals of M-theory cosmology, it seeks to explore other potential ways for dimensional reduction that could lead to cosmological inflation in the early universe.
@@
In contrast to M-theory, its low-energy effective theory, eleven-dimensional supergravity, has been more extensively studied. Comprehensive reviews, such as Freedman:2012zz ; Green:2012pqa , provide a valuable resource for understanding this field. It is believed that eleven-dimensional supergravity serves as the mother theory for lower-dimensional supergravity theories. Numerous papers have delved into various lower-dimensional supergravity theories and their cosmological models DEath:1996ejf ; Yamaguchi:2011kg ; Ferrara:2016ajl ; Lauria:2020rhc ; Ferrara:2013rsa ; Halyo:1996pp ; Kallosh:2010xz , including discussions on no-scale inflation as discussed in references Ellis:2013xoa ; Ellis:2013nxa ; Wu:2024lrw ; Wu:2022kew . However, cosmological inflation models based on the original eleven-dimensional supergravity remain rare. Early discussions have touched upon the inflationary scenario involving the compactification of seven dimensions as $S^{7}$ Hawking:1998ub ; Bremer:1998zp . Other approaches from M-theory to string cosmology such as Calabi-Yau compactification $CY^{3}\times S^{1}$ were studied in Billyard:1999dg ; Cicoli:2008gp .
@@
After inflation, dark matter plays a vital role in cosmic expansion and galaxy formation which can be interpreted as a scalar field, specifically the axion or dilaton field Svrcek:2006yi ; Schive:2014dra ; Marsh:2015daa ; Stadnik:2015kia ; Tenkanen:2019aij ; Appelquist:2024koa ; OHare:2024nmr . The dark scalar field non-minimally coupled to gravity was studied in Refs. Cosme:2018nly ; Alonso-Alvarez:2018tus ; Shi:2019hwe . When considering galactic scales, a scalar field is interpreted as a condensate of ultra-light bosonic dark matter (UBDM) that interacts with general relativity, adhering to the Klein-Gordon equations. In the non-relativistic realm, the equations of motion reduce to the Schrödinger-Poisson system. It is postulated that, at galactic scales, this dark matter forms power-law matter density profiles described by $\rho=\rho_{0}r^{-\alpha}$ Pierobon:2023ozb ; Eggemeier:2024fzs ; Bertschinger:1985pd or the Navarro-Frenk-White matter density profiles Navarro:1995iw ; Primack:1998sa
@@
where $r_{s}$ is the characteristic scale radius. On the other hand, MOdified Newtonian Dynamics (MOND) Milgrom:1983zz ; Milgrom:1983ca ; Bekenstein:1984tv ; Milgrom:1986ib is a possible and promising alternative to the dark matter problem which can explain the flat rotation curves of galaxies and the Tully-Fisher relation $v^{4}\propto M$ Tully:1977fu ; McGaugh:2000sr ; Torres-Flores:2011bna ; Alestas:2021nmi , where $v$ is the rotation speed of a galaxy and $M$ is the total baryonic mass (stars plus gas) in the galaxy. However, it is challenged by the observation of Bullet cluster Clowe:2006eq since the modification to dynamics can not reveal the dark matter distribution. To consider the advantages of the MOND theory, we expect that some hidden matter could induce results that are similar to the implications of MOND.
Another mystery in the dark sector of the universe is dark energy SupernovaSearchTeam:1998fmf ; SupernovaCosmologyProject:1998vns . The dark energy with negative pressure dominates the evolution of our current universe and explains the accelerated expansion of the universe. In the standard cosmological model ($\Lambda$CDM model), the cosmological constant $\Lambda$ as the dark energy contributes about 70% of the energy density of the current universe. Unfortunately, the physical interpretation of the dark energy suffers from the cosmological constant problem or the fine-tuning problem since the value of the energy density of cosmological constant $\rho_{\Lambda}\simeq 10^{-120}m_{p}^{4}$, where $m_{p}\sim 10^{18}\mathrm{GeV}$ is a Planck mass scale Carroll:1991mt ; Martin:2012bt . This problem may be alleviated by the mechanism of small-scale spacetime fluctuations Wang:2019mee or spacetime foam Carlip:2018zsk . In fact, dark energy is not necessary to be a cosmological constant. Recent observations from the Dark Energy Spectroscopic Instrument (DESI), along with other datasets including CMB, PantheonPlus, Union3, and DESY5, support the time-varying dark energy equation of state DESI:2024mwx .
In this study, we use the tree-level eleven-dimensional supergravity, which we will abbreviate as “11D supergravity”, as the starting point. The action of the 11D supergravity is given by Freedman:2012zz
@@
In Section 2, we propose that by compactifying seven dimensions in the form of $\mathcal{M}^{3,1}\times T_{q_{1}}^{2}\times T_{q_{2}}^{2}\times T_{q_{3}}^{2}\times S^{1}$, the action (2) can give rise to the complete evolution of the four-dimensional visible universe, encompassing matter described by the Standard Model of Particle Physics as well as dark matter and dark energy. The four-dimensional extended spacetime, denoted by $\mathcal{M}^{3,1}$, is described using the Friedmann-Robertson-Walker (FRW) metric. The gauge conditions for the gauge parameters $\theta_{MN}$ and $\epsilon$ of the 11D supergravity can be chosen to prevent the extra dimensions from decompactifying as the universe evolves. Through numerical analysis, in Section 3, we will see that a contraction of the 3 tori and an expansion of the eleventh dimension gives rise to cosmological inflation in the four-dimensional FRW universe. The inflaton is interpreted as the conformal vibrations of the extra dimensions. Furthermore, in Section 4, we discussed the quantum effects coming from the compactified extra dimensions, where the inflaton is supposed to be dark matter or dark energy. We will see that the leading order contribution from the four-point matter scattering amplitude results in a Yukawa-like modification to Newton’s law of gravity. Furthermore, this leading order relativistic contribution can explain the Tully-Fisher relation $v^{4}\sim M$ observed in galaxies.
2 Compactification Towards Four Dimensions
@@
The universe we see comprises four extended dimensions: three spatial dimensions and one temporal dimension. In the theory of eleven-dimensional supergravity with all loop-order corrections, the remaining seven dimensions of the theory should be compactified to an extremely small scale. These compactified dimensions are presumably microscopic and their vibrations are influenced by quantum effects. At the beginning of the universe, scattering processes involving gravitons, gravitinos, and 3-forms are prevalent, resulting in a boiling spacetime that can not be perceived as a background of the tree-level eleven-dimensional supergravity (2). In this scenario, the supergravity necessitates the inclusion of an infinite number of loop-order quantum corrections, which lie beyond our current analytical capabilities. Hence, we postulate that the inflationary period commences when the loop-order graviton scattering processes become a negligible contribution to the universe’s total S-matrix, rendering all the loop-order quantum corrections of the graviton field insignificant. In our inflation model, matter fields of four-dimensional spacetime effective action are assumed temporarily inactive during the inflation period, resulting in the graviton and inflaton dominating the universe’s dynamics. This assumption of the inflationary period can be interpreted as selecting suitable gauge conditions for the gauge parameters $\theta_{MN}$ and $\epsilon$ and boundary conditions for the cosmological D3-branes in eleven-dimensional bulk.
2.1 A Special Case
@@
Four extended dimensions within the 11D bulk exhibit the FRW universe:
@@
where $\mathcal{G}_{\mu\nu},\mathcal{G}_{a_{i}b_{i}},i=1,2,3$, are the Einstein tensor for metric $g_{\mu\nu},g_{a_{1}b_{1}},h_{a_{2}b_{2}},\gamma_{a_{3}b_{3}}$ respectively. The energy-momentum tensor, denoted as $T_{AB}^{(A,\psi)}$, is defined for the gravitino $\psi_{A}$ and 3-form $A_{MNP}$ of 11D supergravity. Note that there are gauge parameters $\theta_{MN}$ and $\epsilon$ in 11D supergravity. It is imperative to incorporate this gauge condition when solving the equations of motion for the gravitino and the 3-form. Because the gauge parameter $\theta_{MN}$ has 55 degrees of freedom, we can impose gauge conditions as follows:
1. For a given function $V(\Phi)$ and a symmetric tensor $\mathcal{T}_{\mu\nu}^{\text{(matter)}}$, the equations
$T_{\mu\nu}^{(A,\psi)}=\frac{m_{p(11)}^{9}}{m_{p}^{2}}\mathcal{T_{\mu\nu}^{\text{(matter)}}}-m_{p(11)}^{9}\left[\frac{V(\Phi)}{m_{p}^{2}}+\frac{1}{2}\left(\frac{R_{T_{q_{1}}^{2}}}{\Phi}+R_{T_{q_{2}}^{2}}+\frac{R_{T_{q_{3}}^{2}}}{\Phi^{2}}\right)\right]g_{\mu\nu},$ (22)
expect for $\mu=\nu=0$ case, restrict 9 degrees of freedom.
$G_{a_{i}A}$ $=0,(A\neq a_{i}\text{ and }i=1,2,3)$ (23)
$G_{\mu A}$ $=0,(A\neq\mu)$ (24)
$G_{10,A}$ $=0,(A\neq 10)$ (25)
where the operator $\delta(\cdots)$ is the variation respect to supersymmetry. The supersymmetric transformation law for the metric $G_{AB}$ can be expressed as
$\delta G_{a_{1}b_{1}}$ $=g_{a_{1}b_{1}}\delta\Phi,$ (34)
$\delta G_{a_{2}b_{2}}$ $=0,$ (35)
$\delta G_{a_{3}b_{3}}$ $=2g_{a_{3}b_{3}}\Phi\delta\Phi,$ (36)
$\delta G_{10,10}$ $=-6\Phi^{-7}\delta\Phi.$ (37)
The gauge condition (33) indicates the absence of gravitino, $\psi_{\mu}=0$, in the four-dimensional spacetime. Consequently, supersymmetry of gravity in the four-dimensional spacetime is spontaneous breaking. As there is no strong observable signal from supersymmetric field theory in 4D, we postulate that $\Phi$ has no superpartner, thus assuming $\delta\Phi=0$. Hence, the compactification methods spontaneously break all the $\mathcal{N}=32$ supersymmetries present in the eleven-dimensional bulk, where the classical solutions of (33)(34)(35)(36)(37) are $\psi_{M}=0$. The non-trivial solution to the gauge parameter $\epsilon$ can be derived from the equation
Note that there are trivial solutions $\epsilon=f$, with constant Grassman numbers $f$, when the 3-form takes trivial solution $A_{MNP}=0$. After imposing the gauge conditions (22)(23)(24) and (25)(31)(38), the only two remaining undetermined functions are the scale factor $a(t)$ and the scale function $\Phi(x^{\mu})$ in (15).
In the four-dimensional spacetime, we anticipate that Einstein’s gravity predominantly governs the laws of classical gravity as seen in (16). Moreover, gauge conditions (22) imply that it is possible to obtain a four-dimensional effective action, in which the solutions to the equations of motion are represented by $\Phi(x^{\mu})$ and $a(t)$. While the equations of (22), expect for $\mu=\nu=0$, have already been utilized as gauge conditions to determine the solution of 3-form, these equations do not establish the relation between $\Phi(x^{\mu})$ and $a(t)$. To clarify the relationship between $\Phi(x^{\mu})$ and $a(t)$, we introduce an additional equation, which is the same as (22) but only for the case $\mu=\nu=0$. Subsequently, the solution for $\mathcal{T}^{\text{(matter)}}_{00}[\psi_{A},A_{MNP},a(t),\Phi(x^{\mu}),x^{A}]$ obtained from this equation uniquely defines the Lorentz-invariant symmetric tensor $\mathcal{T}^{\text{(matter)}}_{\mu\nu}[\psi_{A},A_{MNP},a(t),\Phi(x^{\mu}),x^{A}]$, thereby imposing 9 gauge conditions specified in (22). These 9 gauge conditions, along with (23)(24)(25)(31)(38), constrain the solutions $\psi_{A}(a(t),\Phi(x^{\mu}),x^{A})$ and $A_{MNP}(a(t),\Phi(x^{\mu}),x^{A})$ in such a way that they determine $\mathcal{T}^{\text{(matter)}}_{\mu\nu}[a(t),\Phi(x^{\mu}),x^{A}]$ uniquely. This tensor can be derived by the equations of motion for $\Phi(x^{\mu})$ and $a(t)$ in four-dimensional spacetime. The equations (22) for $\mu,\nu=0,1,2,3$ serves to define an effective energy-stress tensor $\mathcal{T}_{\mu\nu}$ as
where $m_{p}$ is the four-dimensional reduced Planck mass and $\mathcal{T}_{\mu\nu}^{\text{(matter)}}$ is the stress tensor of some four-dimensional matter. Hence, in the four-dimensional spacetime, the equation of motion (17) gives Einstein’s equation
$\mathcal{G}_{\mu\nu}=\frac{1}{m_{p}^{2}}\mathcal{T}_{\mu\nu},$ (40)
and the associated effective action is given by
where
The dynamics in the 4D spacetime may be influenced by the coordinates associated with the extra dimensions, as reflected in the energy stress tensor $\mathcal{T}^{\text{(matter)}}_{\mu\nu}$. To address this problem, it is natural to introduce $N$ coincident four-dimensional cosmological branes with Dirichlet boundary conditions, such as $x^{10}=c_{0}$ and $x^{a_{i}}=c_{a_{i}}$ for $i=1,2,3$, where $c_{0}$ and $c_{a_{i}}$ are constants. The gravitino field $\psi_{A\dot{\alpha}\dot{\alpha_{1}}\dot{\alpha_{2}}\dot{\alpha_{3}}}$ gives rise to 56 fermion fields $\psi_{i\dot{\alpha}\dot{\alpha}_{1}\dot{\alpha_{2}}\dot{\alpha_{3}}}$ for $i=4,5,\cdots,10$ on cosmological D3-branes, which contains the fermion fields in the Standard Model of Particle Physics. Additionally, there are 21 spin-1 boson fields $A_{\mu ij}$, 35 spin-0 scalar fields $A_{ijk}$, and 7 spin-0 tensor fields $A_{\mu\nu k}$ for $i,j,k=4,5,\cdots,10$. These fields contain ordinary matter, dark matter, and dark energy. Then, the lagrangian for matter on the cosmological D3-branes can encompass ordinary matter from the Standard Model of Particle Physics $\mathcal{L}_{\text{SM}}$, as well as dark matter $\mathcal{L}_{\text{DM}}$ and dark energy $\mathcal{L}_{\text{DE}}$,
$\mathcal{L}^{\text{(Dirichlet)}}_{\text{matter}}=\mathcal{L}_{\text{SM}}+\mathcal{L}_{\text{DM}}+\mathcal{L}_{\text{DE}},$ (43)
and the associated effective action of the cosmological brane is
$S_{\text{brane}}=\int d^{4}x\sqrt{-g_{4}}\left[\frac{m_{p}^{2}}{2}R_{(4)}-\frac{1}{2}\frac{23m_{p}^{2}}{2\Phi^{2}}\nabla_{\mu}\Phi\nabla^{\mu}\Phi-V(\Phi)+\mathcal{L}^{\text{(Dirichlet)}}_{\text{matter}}\right].$ (44)
This allows us to derive a cosmological inflation model based on a specified driven potential $V(\Phi)$, where the potential can be set as
$V(\Phi)=-\frac{m_{p}^{2}}{2}\left(\frac{R_{T_{q_{1}}^{2}}}{\Phi}+R_{T_{q_{2}}^{2}}+\frac{R_{T_{q_{3}}^{2}}}{\Phi^{2}}\right).$ (45)
@@ 2. Particular block diagonal configuration of 11D bulk metric (15) restricts 46 degrees of freedom
As an example, concentrate on the trivial solutions to the equations of motion for gravitino $\psi_{A}$ and 3-form $A_{MNP}$, where the solutions are $\psi_{A}=0$ and $A_{MNP}=0$. The physical significance of these two solutions is that both the gravitino and the 3-form remain inactive during the entire inflationary period. Consequently, the energy-stress tensor simplifies to $T_{AB}^{(A,\psi)}=0$, leaving the 11D Einstein-Hilbert action as the outcome for action (2). By adopting specific gauge conditions, where $T_{\mu\nu}^{(A,\psi)}=0$ in equation (22), and imposing suitable Dirichlet boundary conditions on the cosmological D3-brane with definition of driven potential (45). we can simplify our calculations by setting $\mathcal{L}^{\text{(Dirichlet)}}_{\text{matter}}=0$. This means the four-dimensional matter fields on the branes are temporarily inactive during the inflation period.
Interestingly, one can deduce a suitable potential $V(\Phi)$ by integrating out the physics arising from various configurations of the extra dimensions within the 11D Einstein-Hilbert action
@@
Due to the block diagonal property of the 11D metric (15), the integral over the 11D bulk can be decomposed into multiple integrals over the 4D spacetime and the compactified extra dimensions, corresponding to each block of the 11D metric
@@
$S_{\text{brane}}=S_{EH}=\int d^{4}x\sqrt{-g_{4}}\left[\frac{m_{p}^{2}}{2}R_{(4)}-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-\frac{4\pi m_{p}^{2}}{A_{0}}\left(1-e^{-\sqrt{\frac{2}{23}}\frac{\phi}{m_{p}}}\right)^{2}\right],$ (53)
where $\phi$ is the inflaton. It is important to note that assigning specific values to the genus $q_{1},q_{2},q_{3}$ and the areas $A_{T_{q_{1}}^{2}},A_{T_{q_{2}}^{2}},A_{T_{q_{3}}^{2}}$ essentially imposes boundary conditions on the metric $G_{AB}$. The trick presented in (47)(48) illustrates one potential method of defining the driven potential $V(\Phi(\phi))$ as
@@
When combining equations (45)(54), the Dirichlet boundary conditions for the cosmological D3-brane can be derived from equations
$R_{T^{2}_{q_{1}}}[c_{a_{1}}]=\frac{16\pi}{A_{0}},\ \ R_{T^{2}_{q_{2}}}[c_{a_{2}}]=-\frac{8\pi}{A_{0}},\ \ R_{T^{2}_{q_{3}}}[c_{a_{3}}]=-\frac{8\pi}{A_{0}}.$ (55)
Therefore, the technique presented in equations (47) and (48) is equivalent to imposing specific Dirichlet boundary conditions. In this approach, the relative stretched sizes of the three tori are represented as $(\Phi,1,\Phi^{2})$. While the specific construction of extra dimensions is given in (13), it should be noted that other configurations of extra dimensions are also conceivable. Before conducting numerical analysis, we intend to explore scenarios involving various sizes combinations of the three tori.
2.2 General Case

### v3

@@ Abstract
We explore compactifications of the form $\mathcal{M}^{3,1}\times T_{q_{1}}^{2}\times T_{q_{2}}^{2}\times T_{q_{3}}^{2}\times S^{1}$ in the framework of 11D supergravity. By imposing suitable gauge conditions and boundary conditions, we find that the FRW universe with four extended spacetime dimensions and seven extremely small compactified spatial dimensions emerges as a solution for the 11D supergravity. These specific compactification methods can produce cosmological inflation that aligns with the observational constraints set by the 2021 BICEP/Keck [1] and Planck 2018 results [2]. In the cosmological inflation models we construct, the inflaton can be interpreted as the conformal vibrations of extra dimensions with a size around $10^{5}$ times the reduced Planck length. Additionally, we offer the theoretical predictions for the mass of the inflaton, and the tree-level Newton’s gravity law between two massive point particles surrounded by a spherically symmetric distribution of the inflaton, which can reproduce the Tully-Fisher relation and explain the flat rotation curves of galaxies.
I Introduction
@@
The eleven-dimensional supergravity is widely believed to be the tree-level low-energy effective theory of M-theory. As an ultraviolet complete theory, M-theory possesses inherent advantages in addressing extremely high-energy scenarios, such as the inflationary period of the early universe. However, our understanding of M-theory remains limited, see [3, 4, 5] for remarkable progress in M-theory. A comprehensive theoretical description is still necessary for understanding physics in the extremely early universe by M-theory. From the perspective of string theory, it is conjectured that by compactifying M-theory on a special manifold of $G_{2}$ holonomy, it could yield the supersymmetric version of Standard Model of Particle Physics and General Relativity [6, 7]. This conjecture has generated some discussions on M-theory cosmology in recent years, particularly regarding $G_{2}$ compactification, see [8, 9, 10] for examples. While this paper does not aim to fully realize the ambitious goals of M-theory cosmology, it seeks to explore other potential ways for dimensional reduction that could lead to cosmological inflation in the early universe.
@@
In contrast to M-theory, its low-energy effective theory, eleven-dimensional supergravity, has been more extensively studied. Comprehensive reviews, such as [11, 12], provide a valuable resource for understanding this field. It is believed that eleven-dimensional supergravity serves as the mother theory for lower-dimensional supergravity theories. Numerous papers have delved into various lower-dimensional supergravity theories and their cosmological models [13, 14, 15, 16, 17, 18, 19], including discussions on no-scale inflation as discussed in references [20, 21, 22, 23]. However, cosmological inflation models based on the original eleven-dimensional supergravity remain rare. Early discussions have touched upon the inflationary scenario involving the compactification of seven dimensions as $S^{7}$ [24, 25]. Other approaches from M-theory to string cosmology such as Calabi-Yau compactification $CY^{3}\times S^{1}$ were studied in [26, 27].
@@
After inflation, dark matter plays a vital role in cosmic expansion and galaxy formation which can be interpreted as a scalar field, specifically the axion or dilaton field [28, 29, 30, 31, 32, 33, 34]. The dark scalar field non-minimally coupled to gravity was studied in Refs. [35, 36, 37]. When considering galactic scales, a scalar field is interpreted as a condensate of ultra-light bosonic dark matter (UBDM) that interacts with general relativity, adhering to the Klein-Gordon equations. In the non-relativistic realm, the equations of motion reduce to the Schrödinger-Poisson system. It is postulated that, at galactic scales, this dark matter forms power-law matter density profiles described by $\rho=\rho_{0}r^{-\alpha}$ [38, 39, 40] or the Navarro-Frenk-White matter density profiles [41, 42]
@@
where $r_{s}$ is the characteristic scale radius. On the other hand, MOdified Newtonian Dynamics (MOND) [43, 44, 45, 46] is a possible and promising alternative to the dark matter problem which can explain the flat rotation curves of galaxies and the Tully-Fisher relation $v^{4}\propto M$ [47, 48, 49, 50], where $v$ is the rotation speed of a galaxy and $M$ is the total baryonic mass (stars plus gas) in the galaxy. However, it is challenged by the observation of Bullet cluster [51] since the modification to dynamics can not reveal the dark matter distribution. To consider the advantages of the MOND theory, we expect that some hidden matter could induce results that are similar to the implications of MOND.
Another mystery in the dark sector of the universe is dark energy [52, 53]. The dark energy with negative pressure dominates the evolution of our current universe and explains the accelerated expansion of the universe. In the standard cosmological model ($\Lambda$CDM model), the cosmological constant $\Lambda$ as the dark energy contributes about 70% of the energy density of the current universe. Unfortunately, the physical interpretation of the dark energy suffers from the cosmological constant problem or the fine-tuning problem since the value of the energy density of cosmological constant $\rho_{\Lambda}\simeq 10^{-120}m_{p}^{4}$, where $m_{p}\sim 10^{18}\mathrm{GeV}$ is a Planck mass scale [54, 55]. This problem may be alleviated by the mechanism of small-scale spacetime fluctuations [56] or spacetime foam [57]. In fact, dark energy is not necessary to be a cosmological constant. Recent observations from the Dark Energy Spectroscopic Instrument (DESI), along with other datasets including CMB, PantheonPlus, Union3, and DESY5, support the time-varying dark energy equation of state [58].
In this study, we use the tree-level eleven-dimensional supergravity, which we will abbreviate as “11D supergravity”, as the starting point. The action of the 11D supergravity is given by [11]
@@
In Section II, we propose that by compactifying seven dimensions in the form of $\mathcal{M}^{3,1}\times T_{q_{1}}^{2}\times T_{q_{2}}^{2}\times T_{q_{3}}^{2}\times S^{1}$, the action (2) can give rise to the complete evolution of the four-dimensional visible universe, encompassing matter described by the Standard Model of Particle Physics as well as dark matter and dark energy. The four-dimensional extended spacetime, denoted by $\mathcal{M}^{3,1}$, is described using the Friedmann-Robertson-Walker (FRW) metric. The gauge conditions for the gauge parameters $\theta_{MN}$ and $\epsilon$ of the 11D supergravity can be chosen to prevent the extra dimensions from decompactifying as the universe evolves. Through numerical analysis, in Section III, we will see that a contraction of the 3 tori and an expansion of the eleventh dimension gives rise to cosmological inflation in the extended FRW universe. The inflaton is interpreted as the conformal vibrations of the extra dimensions. Furthermore, in Section IV, we discussed the quantum effects coming from the compactified extra dimensions, where the inflaton is supposed to be dark matter or dark energy. We will see that the leading order contribution from the four-point matter scattering amplitude results in a Yukawa-like modification to Newton’s law of gravity. Furthermore, this leading order relativistic contribution can explain the Tully-Fisher relation $v^{4}\sim M$ observed in galaxies.
II Compactification Towards Four Dimensions
@@
The universe we see comprises four extended dimensions: three spatial dimensions and one temporal dimension. In the theory of eleven-dimensional supergravity with all loop-order corrections, the remaining seven dimensions of the theory should be compactified to an extremely small scale. These compactified dimensions are presumably microscopic and their vibrations are influenced by quantum effects. At the beginning of the universe, scattering processes involving gravitons, gravitinos, and 3-forms are prevalent, resulting in a boiling 11D spacetime that can not be perceived as a background of the tree-level eleven-dimensional supergravity (2). In this scenario, the supergravity necessitates the inclusion of an infinite number of loop-order quantum corrections, which lie beyond our current analytical capabilities. Hence, we postulate that the inflationary period commences when the loop-order graviton scattering processes become a negligible contribution to the universe’s total S-matrix, rendering all the loop-order quantum corrections of the graviton field insignificant. In our inflation model, matter fields of four-dimensional effective action are assumed temporarily inactive during the inflation period, resulting in the graviton and inflaton dominating the universe’s dynamics. This assumption of the inflationary period can be interpreted as selecting suitable gauge conditions for the gauge parameters $\theta_{MN}$ and $\epsilon$ and boundary conditions for the 11D supergravity.
II.1 A Special Case
@@
Four extended dimensions within the 11D bulk exhibit the FRW metric:
@@
where $\mathcal{G}_{\mu\nu},\mathcal{G}_{a_{i}b_{i}},i=1,2,3$, are the Einstein tensor for metric $g_{\mu\nu},g_{a_{1}b_{1}},h_{a_{2}b_{2}},\gamma_{a_{3}b_{3}}$ respectively. The energy-momentum tensor, denoted as $T_{AB}^{(A,\psi)}$, is defined for the gravitino $\psi_{A}$ and 3-form $A_{MNP}$ of 11D supergravity. Note that there are gauge parameters $\theta_{MN}$ and $\epsilon$ in 11D supergravity. It is imperative to incorporate gauge conditions, see A for more details, when solving the equations of motion for the gravitino and the 3-form. After imposing suitable gauge conditions, the solutions for the gravitino and 3-form can be derived from the equations of motion. Note that the compactification method leads to the spontaneous breaking of all the $\mathcal{N}=32$ supersymmetries of 11D supergravity.
@@
As an example, concentrate on the trivial solutions to the equations of motion for gravitino $\psi_{A}$ and 3-form $A_{MNP}$, where the solutions are $\psi_{A}=0$ and $A_{MNP}=0$ by adopting specific gauge conditions. The physical significance of these two solutions is that both the gravitino and the 3-form remain inactive during the entire inflationary period. Consequently, the energy-stress tensor simplifies to $T_{AB}^{(A,\psi)}=0$, leaving the 11D Einstein-Hilbert action as the outcome for action (2). The gauge conditions for the gravitino and 3-form give the matter coupling term of 4D effective action is $\mathcal{L}_{\text{matter}}=0$. Interestingly, one can deduce a suitable potential $V(\Phi)$ by integrating out the physics arising from various configurations of the extra dimensions within the 11D Einstein-Hilbert action
@@
Due to the block diagonal property of the 11D metric (15), the integral over the 11D bulk can be decomposed into multiple integrals over the 4D extended spacetime and the compactified extra dimensions, corresponding to each block of the 11D metric
@@
$S_{\text{eff}}=S_{EH}=\int d^{4}x\sqrt{-g_{4}}\left[\frac{m_{p}^{2}}{2}R_{(4)}-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-\frac{4\pi m_{p}^{2}}{A_{0}}\left(1-e^{-\sqrt{\frac{2}{23}}\frac{\phi}{m_{p}}}\right)^{2}\right],$ (29)
where $\phi$ is the inflaton. The effective action (29) of the 11D bulk arises from integrating out the physics of seven extra dimensions in the low-energy limit. However, it still discribes an 11D universe with four extended spacetime dimensions and seven extremely small compactified spatial dimensions. This means the extra dimensions need to be sufficiently small and in the low energy limit, the extra dimensions does not contribute to the effective action so that the physics from the extra dimensions are hidden. It is important to note that assigning specific values to the genus $q_{1},q_{2},q_{3}$ and the areas $A_{T_{q_{1}}^{2}},A_{T_{q_{2}}^{2}},A_{T_{q_{3}}^{2}}$ essentially imposes boundary conditions on the metric $G_{AB}$. The trick presented in (23)(24) illustrates one potential method of defining the driven potential $V(\Phi(\phi))$ as
@@
In this approach, the relative stretched sizes of the three tori are represented as $(\Phi,1,\Phi^{2})$. While the specific construction of extra dimensions is given in (13), it should be noted that other configurations of extra dimensions are also conceivable. Before conducting numerical analysis, we intend to explore scenarios involving various sizes combinations of the three tori.
II.2 General Case